OpenStudy (anonymous):

Solve. 3x2 – 7x – 2 = 0

jhonyy9 (jhonyy9):

do you know how you can solving a quadratic ?

jhonyy9 (jhonyy9):

us formula of discriminant

jhonyy9 (jhonyy9):

D = b^2 -4ac

jhonyy9 (jhonyy9):

so you dont know it how ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[3x^{2}-7x-2=0\] Use the formula that jhonyy9 gave to calculate the Discriminant \[D=b^{2}-4ac\] now use the quadratic formula which is \[\frac{ -b \pm \sqrt{D} }{ 2a }\] where a=3 b=-7 and c=-2

OpenStudy (anonymous):

uhm ?

OpenStudy (anonymous):

we use these formula's to solve the quadratic equations - \[ax^{2}+bx+c=0\]

OpenStudy (anonymous):

I have no idea how to do it

OpenStudy (anonymous):

\[x=\frac{-b \pm \sqrt{D} }{ 2a }\]

OpenStudy (anonymous):

just put the values. first to calculate D and then in the above formula.

OpenStudy (anonymous):

is the answer \[x=\frac{ 7\pm \sqrt{73} }{ 6 }\]

OpenStudy (anonymous):

yes :-) ..!!

OpenStudy (anonymous):

yayy